# ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM

### v1

@@ Abstract
Multimodal Large Language Model (MLLM) often suffer from hallucinations. They over-rely on partial cues and generate incorrect responses. Recently, methods like Visual Contrastive Decoding (VCD) and Instruction Contrastive Decoding (ICD) have been proposed to mitigate hallucinations by contrasting predictions from perturbed or negatively prefixed inputs against original outputs. In this work, we uncover that methods like VCD and ICD fundamentally influence the model’s internal attention dynamics. This observation suggests that their effectiveness may not stem merely from surface-level modifications to logits but from deeper shifts in attention distribution. Inspired by this insight, we propose an attention-steerable contrastive decoding framework that directly intervenes in the model’s attention mechanisms to offer a more principled approach to mitigating hallucinations. Specifically, we introduce positive and negative steering as two complementary directions for adapting the model’s internal attention distributions. Rather than passively adjusting logits – as it is commonly done – our method dynamically modulates attention pathways within the contrastive decoding process. This enables selective enhancement or suppression of visual feature contributions in a structured manner. Furthermore, we propose a dynamic selection mechanism to identify text-centric heads – those that predominantly attend to text over visual features – for targeted positive steering, as well as a complementary mechanism to select the most critical visual tokens for negative steering, enabling more effective attention adjustments. Our experiments across multiple MLLM architectures (e.g., LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP) and diverse decoding methods (greedy search, beam search, nucleus sampling) demonstrate that our approach significantly reduces hallucinations and improves the performance on benchmarks such as POPE, CHAIR, and MMHal-Bench, while simultaneously enhancing performance on standard VQA benchmarks, including MMMU, MM-VET, ScienceQA, TextVQA, and GQA.
ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM
Yujun Wang1, Jinhe Bi2, Yunpu Ma2,3, Soeren Pirk1 1Christian-Albrechts-Universität zu Kiel (CAU) 2Ludwig Maximilian University of Munich 3Munich Center for Machine Learning
1 Introduction
@@
Recent advances in large language models (LLMs) (Yang et al., 2024; Touvron et al., 2023; Abdin et al., 2024; Raffel et al., 2023; Brown et al., 2020; Devlin et al., 2019) have led to impressive results in a wide array of natural language processing tasks. Building on these successes, researchers have extended LLMs by visual inputs that enable multi-modal large language models (MLLMs) such as LLaVA (Liu et al., 2023b, 2024b). These MLLMs can handle complex tasks like image captioning Anderson et al. (2018), visual question answering Agrawal et al. (2016), and multimodal dialogue Das et al. (2017). Existing approaches (Dai et al., 2023; Liu et al., 2023b, 2024b; Zhou et al., 2024; Chen et al., 2023a; Alayrac et al., 2022; Bi et al., 2024) show remarkable potential to bridge the gap between vision and language.
@@
Despite these achievements, MLLMs often inherit a critical limitation from LLMs: the tendency to produce hallucinations (Huang et al., 2024b; Bai et al., 2024; Liu et al., 2024a). These hallucinations arise when a model over-relies on partial or misleading cues, generating responses that are incorrect or do not correspond to the provided input.
Figure 1: Impact of VCD and ICD on attention distribution. We conduct an image description task using LLaVA-1.5 on 500 randomly sampled images from the COCO dataset while monitoring the internal attention distribution within the LLM component. We compare the changes in attention under different settings of Visual Contrastive Decoding (VCD), Instruction Contrastive Decoding (ICD), and their combination, relative to the original LLaVA model. The x-axis represents different attention categories: system tokens (sys), visual tokens (vis), textual tokens (text), and output tokens (output). The y-axis indicates the attention difference relative to the original model. VCD (solid blue bars) reduces attention to visual tokens while slightly increasing attention to textual tokens, with a stronger effect as the number of noising steps increases. ICD (hatched bars) exhibits a similar trend, further decreasing visual attention and increasing text attention, where stronger negative prefixes (see text in the legend) result in a more pronounced shift. When combining VCD and ICD (dotted bars), the reduction in visual attention is further amplified, while the focus on textual tokens increases. These findings indicate that the effectiveness of VCD and ICD originates from underlying shifts in the model’s attention distribution rather than solely from the contrastive decoding process.
@@
To mitigate hallucinations, two general strategies have emerged: training-phase interventions and inference-phase interventions. In the training phase, auxiliary supervision Chen et al. (2023b) or reinforcement learning Ben-Kish et al. (2024) can help align model outputs with factual or human-preferred references. However, these approaches require additional data or complex reward modeling, which may be costly or infeasible in certain scenarios. In contrast, inference-phase methods (Zhou et al., 2024; Zhao et al., 2024; Deng et al., 2024; Wang et al., 2024a; Leng et al., 2023) aim to correct or filter erroneous outputs without retraining. Contrastive decoding is particularly appealing as it leverages negatively perturbed or prefixed inputs to steer the model away from hallucinations in a training-free manner. Two notable recent methods for contrastive decoding are Visual Contrastive Decoding (VCD) Leng et al. (2023) that perturbs an input image (e.g., via noising) to generate a “negative result” of logits, which is then subtracted from the original logits to suppress hallucinations, and second, Instruction Contrastive Decoding (ICD) Wang et al. (2024a) that prepends a negative prefix to the prompt (e.g., “You are a confused object detector”) to generate a signal that shifts the model’s predictions away from hallucinated content. Both methods offer a lightweight, yet effective approach to reducing hallucinations. However, upon closer examination, we find that these methods construct contrasting branches through surface-level modifications – either perturbing the image (VCD) or prefixing the prompt (ICD) – without explicitly addressing the underlying cause of hallucinations. Attention steering like OPERA and PAI (Liu et al., 2024c; Huang et al., 2024a) is also a common inference-phase remedy to reduce hallucination. However, PAI introduces the notion of “text inertia” – the tendency of an MLLM to keep generating text-driven content even when the image is removed – but does not articulate why steering the attention matrix is the necessary lever to overcome this inertia.
@@
In our experiments (Fig. 1), we observe that both VCD and ICD consistently cause fundamental shifts in the internal attention distribution: they tend to reduce attention on visual tokens and amplify attention on textual tokens. This insight raises a natural question: why not directly steer the attention mechanism itself? To this end, we propose an Attention-Steerable Contrastive Decoding (ASCD) framework to manipulate attention. Specifically, the attention modification is integrated into a contrastive decoding pipeline to either enhance visual cues or to suppress negative signals. We further develop a dynamic head-selection mechanism to identify “text-centric” heads that disproportionately focus on textual cues, enabling more targeted positive adjustments. In parallel, we introduce a complementary mechanism that restricts negative steering to only the most critical visual tokens, ensuring that suppression is applied solely where necessary to mitigate hallucinations while preserving essential visual details. In summary, our contributions are as follows: (1) We analyze how recent contrastive decoding methods (VCD, ICD) create “negative samples” that fundamentally alter attention; (2) We propose an attention-steerable contrastive decoding method that explicitly modulates attention distributions to offer a more principled way to mitigate hallucinations in the inference phase; (3) We faithfully reproduce VCD and ICD to ensure fair comparison with prior work. Across three representative MLLM backbones (LLaVA-1.5 7B, LLaVA-NeXT 7B, and Phi2-SigLIP), three decoding schemes (greedy, nucleus, and beam search), and three hallucination-focused benchmarks (Rohrbach et al., 2019; Li et al., 2023b; Sun et al., 2023) (POPE, CHAIR, MMHal-Bench), our approach consistently reduces hallucinations and strengthens visual grounding. At the same time, it improves performance on standard VQA benchmarks (Yue et al., 2024; Yu et al., 2024; Lu et al., 2022; Singh et al., 2019; Hudson and Manning, 2019), including MMMU, MM-VET, ScienceQA, TextVQA, and GQA whereas other methods suffer from degraded performance on these benchmarks.
Figure 2: A motivating example of proactive attention steering in a visually ambiguous scenario. (a) shows the conversation context where the “orange” is actually tinted blue. (b) shows how the logits vary based on negative-steering. (c) shows how the logits vary based on positive-steering. (d) illustrates how attention-steerable contrastive decoding, which combines both negative and positive steering in a unified framework, reduce hallucinations and produce perception-driven answers.
2 Related Work
@@
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) have significantly advanced the field of artificial intelligence by integrating vision and language understanding, enabling a wide range of vision-language tasks (Dai et al., 2023; Zhu et al., 2023; Liu et al., 2024b, 2023b; Alayrac et al., 2022; Chen et al., 2023a; Zhou et al., 2024; Zhang et al., 2023a; Rong et al., 2025; Chen et al., 2025a; Yu et al., 2025; Guan et al., 2025; Huang et al., 2024c, 2025; Liu et al., 2025; Zhao et al., 2025b). These models typically follow a two-stage training paradigm: (1) large-scale pretraining on web-scale image-text pairs (Liu et al., 2023b; Li et al., 2023a) to learn cross-modal representations, and (2) visual instruction tuning (Liu et al., 2023a; Bi et al., 2025) on task-specific datasets to enhance multimodal instruction-following capabilities. While this paradigm has led to substantial improvements in vision-language reasoning, MLLMs still face key challenges, such as hallucination – where the model generates content that is inconsistent with the given visual input. (Huang et al., 2024b; Bai et al., 2024; Liu et al., 2024a).
@@
Mitigating Hallucinations in MLLMs. Hallucination in MLLMs is particularly pronounced in open-ended generation tasks, where models may produce content that is not aligned with the provided visual input (Huang et al., 2024a; Jing et al., 2024; Zhang et al., 2023b). Some approaches focus on the mitigation of data bias, scaling-up of vision resolution, and alignment optimization. Lovenia et al. (2024) introduce a technique that mines 95,000 negative samples by replacing original categories, attributes, or quantity information with similar but incorrect alternatives. This fine-grained approach effectively enriches the contrastive signal during training, thereby enhancing the model’s robustness. Chen et al. (2024) propose InternVL, which scales the vision encoder up to 6 billion parameters and processes images with widths ranging from 1,664 to 6,144 pixels. While this method improves visual detail and alignment, it requires significant computational resources for pretraining with large-scale data. Sun et al. (2023) employ Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2022) to align different modalities during training. This optimization strategy leads to a reduction in hallucinations by better integrating visual and textual cues. Bi et al. (2024) propose a representation steering method that effectively mitigates hallucination in multimodal models.
@@
Contrastive Decoding Approaches. Recent work has explored contrastive decoding as an effective, training-free means to mitigate hallucinations Xiao et al. (2025). For instance, Leng et al. (2023) introduced Visual Contrastive Decoding (VCD), which perturbs the input image to generate a negative logit branch that is subtracted from the original predictions, while Wang et al. (2024a) employs a negative prompt to steer outputs away from hallucinated content. Huo et al. (2024) leverages a Context and Text-aware Token Selection (CT2S) strategy to selectively retain the most informative vision tokens in early decoder layers, thereby amplifying beneficial multimodal context and suppressing spurious hallucinations.
3 Preliminaries
@@
Modern MLLMs integrate text and visual inputs based on powerful encoders that enable merging the modalities into a unified representation that is processed by a multi-layer Transformer. While these models enable producing coherent responses, they heavily rely on internal attention mechanisms that dictate how visual and textual cues are combined. As discussed in Section 3.2, subtle variations in these attention distributions can significantly impact the generated output. This observation motivates our approach: by explicitly modulating attention, we aim to enhance visual grounding and mitigate hallucinations.
3.1 MLLM Formulation
@@
Embeddings. A unified input is obtained from encoded image and embedded text:
$\mathbf{Z}=[f_{v}(\mathbf{I});f_{t}(\mathbf{x})].$ (2)
Transformer Architecture. The MLLM processes $\mathbf{Z}$ through $L$ Transformer layers Vaswani et al. (2023):
$\mathbf{H}^{(l)}=\mathrm{TransformerLayer}^{(l)}\!\bigl{(}\mathbf{H}^{(l-1)}\bigr{)},\quad\mathbf{H}^{(0)}=\mathbf{Z}.$ (3)
Output Prediction. The final hidden state $\mathbf{h}_{t}^{(L)}$ is mapped to a probability distribution over the vocabulary:
$p_{\theta}(y_{t}\mid\mathbf{I},\mathbf{x},y_{<t})=\mathrm{Softmax}\!\Bigl{(}\mathbf{h}_{t}^{(L)}W^{P}\Bigr{)},$ (4)
where $W^{P}$ is the output projection matrix.
3.2 Proactive Steering of Attention
In Figure 1, we show how Visual Contrastive Decoding (VCD) and Instruction Contrastive Decoding (ICD) indirectly alter attention distributions. Building on this insight, we now ask: what if we explicitly steer the model’s attention? Figure 2 provides a motivating example, illustrating how actively modulating attention can influence the final logits distribution.
Consider a simple query: “What is the color of the orange here?” The conversation context (Figure 2a) is based on LLaVA-1.5 7B, with a provided image in which the “orange” fruit appears to be tinted blue. We experiment with two distinct attention-steering scenarios: negative-steered logits (Figure 2b) and positive-steered logits (Figure 2c). In each case, we proportionally adjust the visual or textual attention before finalizing the output distribution.
In the negative-steered branch, we reduce attention to visual tokens or boost attention to the textual tokens. As shown in the histogram of logits, the model reduces its reliance on the visual input, causing it to fall back more heavily on the LLM’s inherent priors. As a result, it is more likely to generate answers that align with typical linguistic associations rather than the actual content of the image – insisting the color is “orange”. Conversely, the positive-steered branch increases attention to visual tokens or downgrades textual tokens, making the model more sensitive to the actual (albeit unexpected) color in the image. This leads the model to answer “blue” with higher probability.
In addition to these unidirectional adjustments, we further integrate attention steering into the contrastive decoding framework. Instead of using the original logits directly (as in VCD or ICD), we inject the attention-modulated logits. Mathematically, we redefine the contrastive decoding formulation by replacing the original logits adjustment with a positively steered version:
@@ 3.2 Proactive Steering of Attention
By integrating contrastive decoding with explicit attention manipulation, our attention-steerable contrastive decoding framework (Figure 2d), sharpens the output distribution which enhances the likelihood of the correct response, while reducing the impact of competing distractors.
Figure 3: Distribution of text-centric heads across different models and experiment settings. Each heatmap visualizes how frequently a given head occurs among the most text-focused heads. The panel in the center (a) show the result of LLaVA-1.5 with a generation length of 64 tokens; (b) and (c) show results of the same model with longer generation (512 tokens) and a different image set. Despite these changes, LLaVA-1.5 exhibits minimal JS divergence, which indicates consistent text-centric heads. In contrast, Phi2-SigLIP (d) and LLaVA-NeXT (e) deviate substantially from LLaVA-1.5, revealing model-specific attention biases and higher JS divergence.
3.3 Text-centric Heads
@@
Previously, we have highlighted the impact of adjusting attention. In this section, we discuss which heads in the model are most prone to over-reliance on textual cues. To this end, we conduct an experiment to identify "text-centric" heads, i.e., those with disproportionately high text-to-visual attention ratios, and examine their consistency under different generation conditions and image sets. The experiment setup is detailed in Appendix A.
@@
Results and Observations. Figure 3 shows the resulting heatmaps $F$ for multiple models and generation settings. The panel in the center (a) corresponds to LLaVA-1.5 on $N=500$ images with a generation length of 64 tokens. The two heatmaps at the bottom show results of the same model but with either an increased generation length to 512 tokens (b, bottom left), or using a different set of 500 images (c, bottom right). Despite these changes, the distribution of top text-focus heads remains visually similar, and the small Jensen–Shannon (JS) divergences confirm that these text-centric heads are largely invariant under different sampling conditions for the same model.
@@
In contrast, the Phi2-SigLIP (d, top-left) and LLaVA-NeXT (e, top-right) panels deviate significantly from LLaVA-1.5 even under the same experiment settings, with higher JS divergence. This suggests that each model has its own unique set of heads that consistently favor textual attention over visual cues. However, within a single model, the text-centric heads persist across varied prompts, image sets, and generation lengths.
@@
Implications. The consistent presence of the text-centric heads within the same model indicates that certain heads are inherently prone to focusing on textual signals rather than visual content. In Section 4.2 we describe how this insight can be leveraged to selectively target the problematic heads when applying our positive steering strategy. Rather than uniformly amplifying attention across all heads, we concentrate on those that are most responsible for text-dominant attention, thereby avoiding unnecessary modifications to heads that are well-balanced in their visual-textual focus.
4 Methodology
@@
4.1 Text-centric Head Selection
@@
1:Reference dataset $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$, MLLM with $L$ layers and $H$ heads per layer, Final text-centric head count $\kappa_{\textsc{tch}}$
2:$\mathcal{H}_{\textsc{pos}}$ (set of selected text-centric heads)
3:Initialize a global counter $F\in\mathbb{R}^{L\times H}$ to zeros
4:for $i\leftarrow 1$ to $N$ do
5:     Run MLLM on image $\mathbf{I}_{i}$ (e.g., image description)
6:     for all head $(r,c)$ in layer-head grid do
7:         Compute attention ratio:
$Q_{i}(r,c)=\frac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)}$
8:     end for
9:     Identify top-32 indices of $Q_{i}$ (largest ratios) and store in $\mathcal{I}_{i}$
10:     for all $(r,c)\in\mathcal{I}_{i}$ do
11:         $F(r,c)\leftarrow F(r,c)+1$
12:     end for
13:end for
14:Sort all heads $(r,c)$ in descending order by $F(r,c)$
15:Select top $\kappa_{\textsc{tch}}$ heads:
$\mathcal{H}_{\textsc{pos}}\leftarrow\text{first }\kappa_{\textsc{tch}}\text{ heads in sorted list}$
16:return $\mathcal{H}_{\textsc{pos}}$
@@
As detailed in Algorithm 1, we start by identifying the most text-centric heads using a small reference dataset (e.g., 500 images) for a task (e.g., image description). For each sample, we compute the ratio of textual attention to visual attention (Eq. 6 in Appendix A) and take the top 32 heads with the highest ratio. We accumulate these counts over all samples, then choose the top $\kappa_{\textsc{tch}}$ heads as “text-centric”. This step is motivated by our finding (Section 3.3) that certain heads consistently favor textual content over visual cues.
Algorithm 2 Attention-Steerable Contrastive Decoding

### v2

@@ Abstract
Multimodal large language models (MLLMs) frequently hallucinate by over-committing to spurious visual cues. Prior remedies—Visual and Instruction Contrastive Decoding (VCD, ICD)—mitigate this issue, yet the mechanism remains opaque. We first empirically show that their improvements systematically coincide with redistributions of cross-modal attention. Building on this insight, we propose Attention-Steerable Contrastive Decoding (ASCD), which directly steers the attention scores during decoding. ASCD combines (i) positive steering, which amplifies automatically mined text-centric heads—stable within a model and robust across domains—with (ii) negative steering, which dampens on-the-fly identified critical visual tokens. The method incurs negligible runtime/memory overhead and requires no additional training. Across five MLLM backbones and three decoding schemes, ASCD reduces hallucination on POPE, CHAIR, and MMHal-Bench by up to 38.2% while improving accuracy on standard VQA benchmarks, including MMMU, MM-VET, ScienceQA, TextVQA, and GQA. These results position attention steering as a simple, model-agnostic, and principled route to safer, more faithful multimodal generation.
Introduction
@@
Recent advances in large language models (LLMs) [43, 40, 1, 34, 10, 18] have led to impressive results in a wide array of natural language processing tasks. Building on these successes, researchers have extended LLMs with visual inputs that enable multimodal large language models (MLLMs) such as LLaVA [29, 28]. These MLLMs can handle complex tasks like image captioning [4], visual question answering [2], and multimodal dialogue [16]. Existing approaches [15, 29, 28, 48, 12, 3, 8] show remarkable potential to bridge the gap between vision and language.
@@
Despite these achievements, MLLMs often inherit a critical limitation from LLMs: the tendency to produce hallucinations [20, 6, 30]. These hallucinations arise when a model over-relies on partial or misleading cues, generating responses that are incorrect or do not correspond to the provided input.
Figure 1: Impact of VCD and ICD on attention distribution. On 500 COCO images, we measure how Visual (VCD) and Instruction (ICD) Contrastive Decoding redistribute attention in LLaVA-1.5. Both techniques—and their combination—lower attention on visual tokens (vis) while raising it on textual tokens (text), with stronger perturbations yielding larger shifts. This suggests that the reduction in hallucinations achieved by VCD and ICD is largely attributable to the attendant shifts in cross-modal attention, rather than to the logit-subtraction step alone.
@@
To mitigate hallucinations, two general strategies have emerged: training-phase interventions and inference-phase interventions. In the training phase, auxiliary supervision [14] or reinforcement learning [7] can help align model outputs with factual or human-preferred references. However, these approaches require additional data or complex reward modeling, which may be costly or infeasible in certain scenarios. In contrast, inference-phase methods [48, 47, 17, 41, 24] aim to correct or filter erroneous outputs without retraining. Contrastive decoding is particularly appealing as it leverages negatively perturbed or prefixed inputs to steer the model away from hallucinations in a training-free manner. Two notable recent methods for contrastive decoding are Visual Contrastive Decoding (VCD) [24] that perturbs an input image (e.g., via noising) to generate a “negative result” of logits, which is then subtracted from the original logits to suppress hallucinations, and Instruction Contrastive Decoding (ICD) [41] that prepends a negative prefix to the prompt (e.g., “You are a confused object detector”) to generate a signal that shifts the model’s predictions away from hallucinated content. Both methods offer a lightweight, yet effective approach to reducing hallucinations. However, upon closer examination, we find that these methods construct contrasting branches through surface-level modifications—either perturbing the image (VCD) or prefixing the prompt (ICD)—without explicitly addressing the underlying cause of hallucinations. Attention steering like OPERA, IBD and PAI [31, 50, 19] is also a common inference-phase remedy to reduce hallucinations. However, PAI introduces the notion of “text inertia”—the tendency of an MLLM to keep generating text-driven content even when the image is removed—but does not articulate why steering the attention matrix is the necessary lever to overcome this inertia.
@@
To motivate our approach, we first quantify how VCD and ICD reshape a model’s internal attention. As evidenced by Fig. 1, both techniques produce a systematic reallocation of attention from visual tokens to textual tokens. This insight raises a natural question: why not directly steer the attention mechanism itself? To this end, we propose an Attention-Steerable Contrastive Decoding (ASCD) framework to manipulate attention. Specifically, the attention modification is integrated into a contrastive decoding pipeline to both enhance visual cues and suppress negative signals. We further develop a dynamic head-selection mechanism to identify “text-centric” heads that disproportionately focus on textual cues, enabling more targeted positive adjustments. In parallel, we introduce a complementary mechanism that restricts negative steering to only the most critical visual tokens, ensuring that suppression is applied solely where necessary to mitigate hallucinations while preserving essential visual details. In summary, our contributions are as follows: (1) We analyze how recent contrastive decoding methods (VCD, ICD) create “negative samples” that fundamentally alter attention; (2) We propose an attention-steerable contrastive decoding method that explicitly modulates attention distributions to offer a more principled way to mitigate hallucinations in the inference phase; (3) We faithfully reproduce VCD and ICD to ensure fair comparison with prior work. Across five representative MLLM backbones (LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, LLaVA-1.5 13B and Qwen2.5-VL-Instruct), three decoding schemes (greedy, nucleus, and beam search), and three hallucination-focused benchmarks [35, 26, 39] (POPE, CHAIR, MMHal-Bench), our approach consistently reduces hallucinations and strengthens visual grounding. At the same time, it improves performance on standard VQA benchmarks [45, 44, 33, 37, 21], including MMMU, MM-VET, ScienceQA, TextVQA, and GQA whereas other methods suffer from degraded performance on these benchmarks.
Figure 2: A motivating example of proactive attention steering in a visually ambiguous scenario. Top: Conversation context in which the “orange” appears blue-tinted. Middle: Effects of negative steering (decrease vision attention / increase text attention) and positive steering (increase vision attention / decrease text attention); ASCD contrasts the two steered logits to suppress hallucination and produce the perception-consistent answer. Bottom: Color-token logits change with the steering strength for visual and textual attention, corresponding to the steering above.
Related Work
@@
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) have significantly advanced the field of artificial intelligence by integrating vision and language understanding, enabling a wide range of vision-language tasks [15, 49, 28, 29, 5, 48, 36, 11]. These models typically follow a two-stage training paradigm: (1) large-scale pretraining on web-scale image-text pairs [29, 25] to learn cross-modal representations, and (2) visual instruction tuning [27, 9] on task-specific datasets to enhance multimodal instruction-following capabilities. While this paradigm has led to substantial improvements in vision-language reasoning, MLLMs still face key challenges, such as hallucination – where the model generates content that is inconsistent with the given visual input. [20, 6, 30].
@@
Mitigating Hallucinations in MLLMs. Hallucinations in MLLMs are particularly pronounced in open-ended generation tasks, where models may produce content that is not aligned with the provided visual input [19, 23, 46]. Some approaches focus on the mitigation of data bias, scaling-up of vision resolution, and alignment optimization. Lovenia et al. [32] introduce a technique that mines 95,000 negative samples by replacing original categories, attributes, or quantity information with similar but incorrect alternatives. This fine-grained approach effectively enriches the contrastive signal during training, thereby enhancing the model’s robustness. Chen et al. [13] propose InternVL, which scales the vision encoder up to 6 billion parameters and processes images with widths ranging from 1,664 to 6,144 pixels. While this method improves visual detail and alignment, it requires significant computational resources for pretraining with large-scale data. Sun et al. [39] employ Reinforcement Learning from Human Feedback (RLHF) [38] to align different modalities during training. This optimization strategy leads to a reduction in hallucinations by better integrating visual and textual cues. Bi et al. [8] propose a representation steering method that effectively mitigates hallucination in multimodal models.
@@
Contrastive Decoding Approaches. Recent work has explored contrastive decoding as an effective, training-free means to mitigate hallucinations [42]. For instance, Leng et al. [24] introduced Visual Contrastive Decoding (VCD), which perturbs the input image to generate a negative logit branch that is subtracted from the original predictions, while Wang et al. [41] employs a negative prompt to steer outputs away from hallucinated content. Huo et al. [22] leverage a Context and Text-aware Token Selection (CT2S) strategy to selectively retain the most informative vision tokens in early decoder layers, thereby amplifying beneficial multimodal context and suppressing spurious hallucinations.
Preliminaries
@@
Modern MLLMs integrate text and visual inputs based on powerful encoders that enable the merging of the modalities into a unified representation that is processed by a multi-layer Transformer. While these models enable the production of coherent responses, they heavily rely on internal attention mechanisms that dictate how visual and textual cues are combined. As discussed in previous section, subtle variations in these attention distributions can significantly impact the generated output. This observation motivates our approach: by explicitly modulating attention, we aim to enhance visual grounding and mitigate hallucinations.
MLLM Formulation
@@
Transformer Backbone.
The input is processed by $L$ Transformer blocks, and each block contains $H$ attention heads. We denote the unnormalized attention score matrix of the head $h$ in layer $l$ by $\mathbf{A}^{(l)}_{h}$.
Proactive Steering of Attention
In Figure 1, we show how visual contrastive decoding (VCD) and instruction contrastive decoding (ICD) indirectly alter attention distributions. Building on this insight, we now ask: what if we explicitly steer the model’s attention? Figure 2 provides a motivating example, illustrating how actively modulating attention can influence the final logit distribution.
Consider a simple query: “What is the color of the orange here?” The conversation context (Figure 2) is based on LLaVA-1.5 7B, with a provided image in which the “orange” fruit appears to be tinted blue. We experiment with two distinct attention-steering scenarios: negative-steered logits and positive-steered logits. In each case, we proportionally adjust the visual or textual attention before finalizing the output distribution.
In the negative-steered branch, we reduce attention to visual tokens or boost attention to the textual tokens. As shown in the histogram of logits, the model reduces its reliance on the visual input, causing it to fall back more heavily on the LLM’s inherent priors. As a result, it is more likely to generate answers that align with typical linguistic associations rather than the actual content of the image—insisting that the color is “orange”. Conversely, the positive-steered branch increases attention to visual tokens or downgrades textual tokens, making the model more sensitive to the actual (albeit unexpected) color in the image. This leads the model to answer “blue” with higher probability.
In addition to these unidirectional adjustments, we further integrate attention steering into the contrastive decoding framework. Instead of using the original logits for the positive branch directly (as in VCD or ICD), we inject the attention-modulated logits. Mathematically, we redefine the contrastive decoding formulation by replacing the original logits adjustment with a positively steered version:
@@
By integrating contrastive decoding with explicit attention manipulation, our attention-steerable contrastive decoding framework (Figure 2 right) sharpens the output distribution, enhances the likelihood of the correct response, while reducing the impact of competing distractors.
Methodology
Figure 3: The stability of text-centric head distribution. Each heatmap visualizes how frequently a given head occurs among the most text-focused heads. LLaVA-1.5(a) remains stable across generation length(b) and image set(c), whereas Phi2-SigLIP(d) and LLaVA-NeXT(e) shift markedly.
Text-centric Heads are Model-specific
@@
Having established the impact of attention adjustments, we now discuss which heads in the model are most prone to over-reliance on textual cues. Here, we conduct an experiment to identify "text-centric" heads, i.e., those with disproportionately high text-to-visual attention ratios, and examine their consistency under different generation conditions and image sets. The experimental setup is detailed in the supplementary material (Appendix, Text-Centric Heads Experiment Settings).
@@
Results and Observations. Figure 3 shows the resulting heatmaps $F$ for multiple models and generation settings. The panel in the center (a) corresponds to LLaVA-1.5 on $N=500$ images with a generation length of 64 tokens. The two heatmaps at the bottom show results of the same model but with either an increased generation length to 512 tokens (b, bottom left) or using a different set of 500 images (c, bottom right). Despite these changes, the distribution of top text-focused heads remains visually similar, and the small Jensen–Shannon (JS) divergences confirm that these text-centric heads are largely invariant under different sampling conditions for the same model.
@@
In contrast, the Phi2-SigLIP (d, top-left) and LLaVA-NeXT (e, top-right) panels deviate significantly from LLaVA-1.5 even under the same experimental settings, with higher JS divergence. This suggests that each model has its own unique set of heads that consistently favor textual attention over visual cues. However, within a single model, the text-centric heads persist across varied prompts, image sets, and generation lengths.
@@
Implications. The consistent presence of the text-centric heads within the same model indicates that certain heads are inherently prone to focusing on textual signals rather than visual content. In the next subsection, we describe how this insight can be leveraged to selectively target the problematic heads when applying our positive steering strategy.
Robustness across Data Domains. To further assess the robustness of model-specific text-centric heads, we repeat the profiling on extremely out-of-domain X-ray data. Details are provided in the supplementary material (“Extended Analysis of Text-centric Heads on Medical Data Domain”). The resulting heatmaps and Jensen–Shannon divergences show that, within each model, the same subset of heads remains text-centric despite the shift from COCO photographs to chest X-rays, indicating strong domain robustness.
@@
Text-centric Head Selection
Input: Reference image set $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$; MLLM with $L$ layers and $H$ heads per layer; desired text-centric head count $\kappa_{\textsc{tch}}$
Output: $\mathcal{H}_{\textsc{pos}}$ // Selected text-centric heads
Initialize Global Statistics:
Initialize counter tensor $F\leftarrow\mathbf{0}^{L\times H}$ // Vote accumulator
[2pt]
Vote Accumulation Over Reference Set:
foreach $\mathbf{I}_{i}\in\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$ do // Iterate over images
   Run MLLM on $\mathbf{I}_{i}$ to obtain cross-modal attentions
 foreach $(r,c)\in\{1,\dots,L\}\times\{1,\dots,H\}$ do // Loop over all heads
    $Q_{i}(r,c)\leftarrow\dfrac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)}$
 $\mathcal{I}_{i}\leftarrow$ indices of the top-32 $Q_{i}$ values
 foreach $(r,c)\in\mathcal{I}_{i}$ do
    $F(r,c)\leftarrow F(r,c)+1$
Head Selection:
Sort heads $(r,c)$ by $F(r,c)$ in descending order
$\mathcal{H}_{\textsc{pos}}\leftarrow$ top $\kappa_{\textsc{tch}}$ heads
return $\mathcal{H}_{\textsc{pos}}$
@@
As detailed in Algorithm 1, we start by identifying the most text-centric heads using a small reference dataset (e.g., 500 images) for a task (e.g., image description). For each sample, we compute the ratio of textual attention to visual attention and take the top 32 heads with the highest ratio. We accumulate these counts over all samples, then choose the top $\kappa_{\textsc{tch}}$ heads as “text-centric”. This step is motivated by our previous finding that certain heads consistently favor textual content over visual cues.